OpenStudy (anonymous):

Solve: logx+5 36 = 2

OpenStudy (anonymous):

logx = 2-536 logx=-534 x=10^(-534)

OpenStudy (anonymous):

The base is x+5

OpenStudy (experimentx):

36 = (x+5)^2 or -6 = x+5 => x= -11 or +6 = x+5 => x= -1

OpenStudy (lgbasallote):

x= 6-5 is not equal to x=-1 just saying :)

OpenStudy (experimentx):

yep